According to a Pew Research Center study, in May 2011, 31% of all American adults had a smart phone (one which the user can use to read email and surf the Internet). A communications professor at a university believes this percentage is higher among community college students. She selects 352 community college students at random and finds that 129 of them have a smart phone. In testing the hypotheses:

H0: p = 0.31 versus

Ha: p > 0.31,

she calculates the test statistic as z = 2.2911.

Then the p‑value = .

(Please round your answer to four decimal places.)
Answer #1

solution:

In testing the hypotheses:

this is right tail test

H0: p = 0.31 versus

Ha: p > 0.31

test statistic as z = 2.2911.

p(z>2.2911)=1-p(z<2.2911)=1-0.9890=0.0110

p value =0.0110